# Multi-Agent Autonomous Driving Systems with Large Language Models: A Survey of Recent Advances

### v1

@@ Abstract
Autonomous Driving Systems (ADSs) are revolutionizing transportation by reducing human intervention, improving operational efficiency, and enhancing safety. Large Language Models (LLMs), known for their exceptional planning and reasoning capabilities, have been integrated into ADSs to assist with driving decision-making. However, LLM-based single-agent ADSs face three major challenges: limited perception, insufficient collaboration, and high computational demands. To address these issues, recent advancements in LLM-based multi-agent ADSs have focused on improving inter-agent communication and cooperation. This paper provides a frontier survey of LLM-based multi-agent ADSs. We begin with a background introduction to related concepts, followed by a categorization of existing LLM-based approaches based on different agent interaction modes. We then discuss agent-human interactions in scenarios where LLM-based agents engage with humans. Finally, we summarize key applications, datasets, and challenges in this field to support future researchWe provide an open-source library for future study via the following link: https://anonymous.4open.science/r/LLM-based_Multi-agent_ADS-3A5C/README.md.
@@
A Survey of Recent Advances
Yaozu Wu1,, Dongyuan Li1,, Yankai Chen2,3,, Renhe Jiang1,, Henry Peng Zou3, Liancheng Fang3, Zhen Wang1, Philip S. Yu3 1The University of Tokyo, 2Cornell University, 3University of Illinois Chicago {yaozuwu279,zhenwangr}@gmail.com, {lidy,jiangrh}@csis.u-tokyo.ac.jp, yankaichen@acm.org, {pzou3,lfang87,psyu}@uic.edu
@@ 1 Introduction
Autonomous driving systems (ADSs) are redefining driving behaviors, reshaping global transportation networks, and driving a technological revolution Yurtsever et al. (2020). Traditional ADSs primarily rely on data-driven approaches (as detailed in Appendix A.1), often focusing on system development while overlooking dynamic interactions with the environment. To enhance engagement with diverse and complex driving scenarios, agentic roles have been incorporated into ADSs Durante et al. (2024) using methods such as reinforcement learning Zhang et al. (2024b) and active learning Lu et al. (2024). Despite notable progress, these methods struggle with “long-tail" scenarios, where rare but critical driving situations—such as sudden obstacles—pose significant challenges to model performance. Furthermore, their “black-box" nature limits interpretability, making their decisions difficult to trust.
LLM-based single-agent ADSs help overcome the limitations of data-driven methods Wang et al. (2024a). Pre-trained on vast, multi-domain datasets, LLMs excel in knowledge transfer and generalization Achiam et al. (2023), enabling strong performance in traffic scenarios under zero-shot settings, thus addressing the long-tail issue Yang et al. (2023). Moreover, techniques such as Reinforcement Learning from Human Feedback (RLHF) and Chain-of-Thought (CoT) Zhao et al. (2023), enhance language-based interaction and logical reasoning, allowing LLMs to make human-like, real-time decisions while providing interpretable and trustworthy feedback across various driving conditions. For instance, Drive-Like-a-Human Fu et al. (2024) builds a closed-loop system comprising environment, agent, memory, and expert modules. The agent interacts with the environment, reflects on expert feedback, and ultimately accumulates experience. DiLu Wen et al. (2024) replaces human experts with a reflection module and integrates an LLM-based reasoning engine to enable continuous decision-making. Agent-Driver Mao et al. (2024) designs a tool library to collect environmental data and uses LLMs’ cognitive memory and reasoning to improve planning.
@@
However, as shown in Figure 1, researchers have identified three critical limitations of LLM-based single-agent ADSs in complex traffic environments: ❶ Limited Perception: LLMs can only respond to sensor inputs and lack predictive and generalization capabilities. As a result, LLM-based single-agent ADSs cannot complement incomplete sensor information and thus miss critical information in driving scenarios, such as pedestrians or vehicles hidden in complex intersection environments Hu et al. (2024c). ❷ Insufficient Collaboration: A single LLM-based agent cannot coordinate with other vehicles or infrastructure, leading to suboptimal performance in scenarios requiring multi-agent interactions, such as lane merging or navigating roundabouts Hu et al. (2021). ❸ High Computational Demands: With billions of parameters in LLMs, these methods demand substantial independent computational resources, making real-time deployment challenging, particularly in resource-limited in-vehicle systems Cui et al. (2023).
@@
To address these limitations, LLM-based multi-agent ADSs enable distinct agents to communicate and collaborate, enhancing safety and performance. First, LLMs enhance contextual awareness by allowing agents to share data, extend their perceptual range, and enhance the detection of occluded objects in complex environments Hu et al. (2024c). Second, real-time coordination between LLM-based agents mitigates insufficient collaboration, enabling joint decision-making in scenarios such as lane merging and roundabout navigation, ultimately leading to safer and more efficient driving operations Hu et al. (2021). Third, LLMs optimize computational efficiency by distributing tasks among agents, reducing individual workloads, and enabling real-time processing in resource-limited systems Cui et al. (2023).
@@
As LLM capabilities continue to advance, they are playing an increasingly significant role in ADS as intelligent driving assistants. Several reviews have focused on two primary aspects: i) the integration of LLMs into data-driven methods Yang et al. (2023); Li et al. (2023) and ii) the applications of specific LLM types, such as vision-based Zhou et al. (2024b) and multimodal-based Fourati et al. (2024); Cui et al. (2024c) models in ADSs. However, no comprehensive survey has systematically examined the emerging field of LLM-based multi-agent ADSs. This gap motivates us to provide a thorough review that consolidates existing knowledge and offers insights to guide future research and the development of advanced ADSs.
@@
In this study, we present a comprehensive survey of LLM-based multi-agent systems. Specifically, Section 2 introduces the core concepts of LLM-based multi-agent ADSs, including agent environments and profiles, inter-agent interaction mechanisms, and agent-human interactions. Section 3 provides a structured review of the state-of-the-art in multi-agent ADS, categorizing existing studies into three key interaction types: multi-vehicle interaction, vehicle-infrastructure interaction, and vehicle-assistant interaction. As agent capabilities continue to grow, human-vehicle co-driving is becoming the dominant autonomous driving paradigm, with human involvement playing an increasingly vital role. Humans collaborate with agents by providing guidance or supervising their behavior. Therefore, we consider humans as special virtual agents and examine human-agent interactions in Section 4. Section 5 explores various applications, while Section 6 compiles a comprehensive collection of public datasets and open-source resources. Section 7 discusses existing challenges and future research directions and Section 8 concludes the study.
@@ 2.2.1 Agent Environment and Profile
  • Physical environment. It represents the real-world setting where driver agents gather information using various sensors, such as cameras and LiDAR, and interact with other traffic participants. However, due to the high cost of vehicles and strict regulations on public roads, collecting large amounts of data in real world is impractical.
  • Simulation environment. As a viable alternative, the simulation environment provides a simulated setting constructed by humans. It can accurately model specific conditions without incurring the high costs and complexities associated with real-world data collection, allowing agents to freely test actions and strategies across a variety of scenarios Dosovitskiy et al. (2017).
@@
  • Pre-defined methods. In these cases, system designers explicitly define agent profiles based on prior knowledge and the analysis of complex scenarios Chen et al. (2024a). Each agent has unique attributes and behavior patterns that can be adjusted based on the scenario. In driving environments, the objectives of ADS require the collaboration of vehicle agents, infrastructure agents, and drivers. In particular, ❶ Vehicle agents denote various types of autonomous vehicles, traveling according to preset routes and traffic rules, while communicating and collaborating with other vehicles and driver agents. ❷ Infrastructure agents, e.g., traffic lights, road condition monitors, and parking facilities, provide real-time traffic information and instructions, influencing the behavior of driver and vehicle agents.
  • Model-generated methods. These approaches create agent profiles using advanced LLMs based on the interaction context and the goals that need to be accomplished Zhou et al. (2024c).
  • Data-derived Profile. They design agent profiles based on pre-existing datasets Guo et al. (2024).
@@ 2.2.2 LLM-based Multi-Agent Interaction
In LLM-based multi-agent ADS, effective information exchange and action coordination between agents are essential to improve collective intelligence and solve complex traffic scenarios. Agent interactions are influenced by both the interaction mode and the underlying interaction structure.
  • The interaction mode of LLM-based multi-agent ADS can be classified as: cooperative, competitive, and debate mode. ❶ In cooperative mode, agents work together to achieve shared objectives by exchanging information Chen et al. (2024d); Jin et al. (2024). ❷ In competitive mode, agents strive to accomplish their individual goals and compete with others Yao et al. (2024). ❸ The Debate mode enables agents to debate with each other, propose their own solutions, criticize the solutions of other agents, and collaboratively identify optimal strategies Liang et al. (2024).
@@
    Figure 3: Different interaction modes and interaction structures.
    Specifically, ❶ the centralized interaction structures defines a central agent or a group of central agents to manage interactions among all agents Zhou et al. (2024c). ❷ The decentralized interaction structure allows for direct communication between agents, with all agents being equal to each other Hu et al. (2024b). ❸ Hierarchical structures focus on interactions within a layer or with adjacent layers Ohmer et al. (2022). ❹ The shared memory interaction structure maintains a shared message pool, allowing agents to send and extract the necessary information Jiang et al. (2024). We provide a more detailed introduction to LLM-based multi-agent ADSs based on their interaction structures and modes in Section 3.
@@ 2.2.3 LLM-based Agent-Human Interaction
Recent studies have shown that human-machine co-driving systems leverage LLMs to improve agent-human interactions, enabling autonomous vehicles to communicate and collaborate seamlessly with human drivers through natural language Feng et al. (2024). This capability allows vehicles to better understand and respond to human intent, provide context-aware responses, enhance driving safety and comfort, and offer personalized recommendations based on driver preferences. Furthermore, humans play a crucial role in guiding and supervising agent behavior, enhancing the agents’ capabilities while ensuring safety and compliance with legal standards. We explore the role of humans as special virtual agents in LLM-based multi-agent ADS and examine the intricate dynamics of agent-human interactions in Section 4.
@@ 3 LLM-based Multi-Agent Interaction
{forest}
[…]
Assistance-Tools, fill=myMintpink, font=, text width=5.5em [ChatSim Wei et al. (2024), HumanSim Zhou et al. (2024a), EC-Drive Chen et al. (2024a),
CAV-LLM-Driving-Assistant Tang et al. (2024), LLM-Edge-Device-Driving Huang et al. (2024b),
ALGPT Zhou et al. (2024c), fill=myMintpink, font=, text width=25.8em] ] [ Collaborative
Decision-Making, fill=myMintpink, font=, text width=5.5em [LanguageMPC Sha et al. (2023), CoMAL Yao et al. (2024), AgentsCoDriver Hu et al. (2024a),
AgentsCoMerge Hu et al. (2024b), CoDrivingLLM Fang et al. (2024), KoMA Jiang et al. (2024),
SurrealDriver Jin et al. (2024), fill=myMintpink, font=, text width=25.8em] ] [ Collaborative
Perception, fill=myMintpink, font=, text width=5.5em [Complement-Vehicle’s-FOV Dona et al. (2024), V-HOI MLCR Zhang et al. (2024a), fill=myMintpink, font=, text width=25.8em] ] ] [ Human-Agent
Interaction, fill=lime!5, font=, text width=5em [ Partnership
Paradigm, fill=lime!5, font=, text width=5.5em [Drive-as-You-Speak Cui et al. (2024a), Reason-and-React Cui et al. (2024b),
DriVLMe Huang et al. (2024a), AccidentGPT Wang et al. (2024b),
ConnectGPT Tong and Solmaz (2024), fill=lime!5, font=, text width=25.8em ] ] [ Instructor
Paradigm, fill=lime!5, font=, text width=5.5em [Co-Pilot Wang et al. (2023), PPE Ma et al. (2024), , fill=lime!5, font=, text width=25.8em ] ] ] [ Multi-Agent
Interaction, fill=blue!5, font=, text width=5em [ Vehicle-Assistant
Interaction, fill=blue!5, font=, text width=5.5em [ChatSim Wei et al. (2024), ALGPT Zhou et al. (2024c), AD-H Zhang et al. (2024c),
SurrealDriver Jin et al. (2024), LDPD Liu et al. (2024a), V-HOI MLCR Zhang et al. (2024a), fill=blue!5, font=, text width=25.8em] ] [ Vehicle-Infrastructure
Interaction, fill=blue!5, font=, text width=5.5em [CAV-LLM-Driving-Assistant Tang et al. (2024), EC-Drive Chen et al. (2024a), fill=blue!5, font=, text width=25.8em] ] [ Multi-Vehicle
Interaction, fill=blue!5, font=, text width=5.5em [LanguageMPC Sha et al. (2023), AgentsCoDriver Hu et al. (2024a), KoMA Jiang et al. (2024),
AgentsCoMerge Hu et al. (2024b), CoDrivingLLM Fang et al. (2024),CoMAL Yao et al. (2024),
Complement-Vehicle’s-FOV Dona et al. (2024), fill=blue!5, font=, text width=25.8em] ] ] ]
@@ 3.1 Multi-Vehicle Interaction
Multi-vehicle interactions involve multiple autonomous vehicles powered by LLMs exchanging real-time information, such as locations, speeds, sensor data, and intended trajectories. By sharing partial observations of the environment or negotiating maneuvers, multiple vehicles overcome the inherent limitations of single-agent ADS, such as restricted perception and lack of collaboration. Typically, these interactions operate in a cooperative mode. LanguageMPC Sha et al. (2023) employs a centralized structure, where a central agent acts as the “brain” of the fleet, providing coordination and control commands to each vehicle agent. In contrast, other decentralized approaches Fang et al. (2024); Dona et al. (2024) treat all agents equally, allowing direct communication between multiple agents. For instance, AgentsCoDriver Hu et al. (2024a) designs a communication module that generates messages for inter-agent communication when the agent deems it necessary. AgentsCoMerge Hu et al. (2024b) and CoDrivingLLM Fang et al. (2024) incorporate agent communication into the reasoning process, facilitating intention sharing and negotiation before decision-making. Additionally, KoMA Jiang et al. (2024) and CoMAL Yao et al. (2024) build a shared memory pool, allowing agents to send and retrieve the necessary information to facilitate interaction between agents.

### v2

@@ Abstract
Autonomous Driving Systems (ADSs) are revolutionizing transportation by reducing human intervention, improving operational efficiency, and enhancing safety. Large Language Models (LLMs) have been integrated into ADSs to support high-level decision-making through their powerful reasoning, instruction-following, and communication abilities. However, LLM-based single-agent ADSs face three major challenges: limited perception, insufficient collaboration, and high computational demands. To address these issues, recent advances in LLM-based multi-agent ADSs leverage language-driven communication and coordination to enhance inter-agent collaboration. This paper provides a frontier survey of this emerging intersection between NLP and multi-agent ADSs. We begin with a background introduction to related concepts, followed by a categorization of existing LLM-based methods based on different agent interaction modes. We then discuss agent-human interactions in scenarios where LLM-based agents engage with humans. Finally, we summarize key applications, datasets, and challenges to support future researchhttps://github.com/Yaozuwu/LLM-based_Multi-agent_ADS.
@@
A Survey of Recent Advances, Resources, and Future Directions
[…]
@@ 1 Introduction
Autonomous driving systems (ADSs) are redefining driving behaviors, reshaping global transportation networks, and driving a technological revolution Yurtsever et al. (2020). Traditional ADSs primarily rely on data-driven approaches (as detailed in Appendix A), focusing on system development while overlooking dynamic interactions with the environment. To enhance engagement with diverse and complex driving scenarios, agentic roles have been incorporated into ADSs Durante et al. (2024) using methods like reinforcement learning Zhang et al. (2024b) and active learning Lu et al. (2024). Despite notable progress, these methods struggle with “long-tail” scenarios, where rare but critical driving situations, such as sudden obstacles, pose significant challenges to model performance. Furthermore, their “black-box” nature limits interpretability, making their decisions difficult to trust.
LLM-based single-agent ADSs help overcome the limitations of data-driven methods Wang et al. (2024a). Pre-trained on vast, multi-domain datasets, LLMs excel in knowledge transfer and generalization Achiam et al. (2023), enabling strong performance in traffic scenarios under zero-shot settings, thus addressing the long-tail issue Yang et al. (2023). Moreover, techniques such as Reinforcement Learning from Human Feedback (RLHF) and Chain-of-Thought (CoT) Zhao et al. (2023), enhance language-based interaction and logical reasoning, allowing LLMs to make human-like, real-time decisions while providing interpretable and trustworthy feedback across various driving conditions. For instance, Drive-Like-a-Human Fu et al. (2024) builds a closed-loop system comprising environment, agent, memory, and expert modules. The agent interacts with the environment, reflects on expert feedback, and ultimately accumulates experience. For example, DiLu Wen et al. (2024) replaces human experts with a reflection module and integrates an LLM-based reasoning engine to enable continuous decision-making. Agent-Driver Mao et al. (2024) designs a tool library to collect environmental data and uses LLMs’ cognitive memory and reasoning to improve planning.
@@
However, as shown in Figure 1, researchers have identified three critical limitations of LLM-based single-agent ADSs in complex traffic environments: ❶ Limited Perception: LLMs can only respond to sensor inputs and lack predictive and generalization capabilities. As a result, LLM-based single-agent ADSs cannot complement incomplete sensor information and thus miss critical information in driving scenarios, such as pedestrians or vehicles hidden in complex intersection environments Hu et al. (2024c). ❷ Insufficient Collaboration: A single LLM-based agent cannot coordinate with other vehicles or infrastructure, leading to suboptimal performance in scenarios requiring multi-agent interactions, such as merging of lanes or navigate roundabouts Hu et al. (2021). ❸ High Computational Demands: With billions of parameters in LLMs, these methods require substantial independent computational resources, making real-time deployment challenging, particularly in resource-limited in-vehicle systems Cui et al. (2023).
@@
To address these limitations, LLM-based multi-agent ADSs enable distinct agents to communicate and collaborate, improving safety and performance. First, LLMs enhance contextual awareness by allowing agents to share data, extend their perceptual range, and enhance the detection of occluded objects in complex environments Hu et al. (2024c). Second, real-time coordination among LLM-based agents mitigates insufficient collaboration, enabling joint decisions in tasks like lane merging and roundabout navigation, ultimately leading to safer and more efficient driving operations Hu et al. (2021). Third, LLMs optimize computational efficiency by distributing tasks across agents, reducing individual load and enabling real-time processing in resource-limited systems.
@@
As LLM capabilities continue to advance, they are playing an increasingly significant role in ADS as intelligent driving assistants. Several reviews have focused on two primary aspects: i) the integration of LLMs in data-driven methods Yang et al. (2023); Li et al. (2023) and ii) the applications of specific LLM types, such as vision-based Zhou et al. (2024b) and multimodal-based Fourati et al. (2024); Cui et al. (2024c) models in ADS. However, no comprehensive survey has systematically examined the emerging field of LLM-based multi-agent ADSs. This gap motivates us to provide a comprehensive review that consolidates existing knowledge and offers insights to guide future research and the development of advanced ADSs.
@@
In this study, we present a comprehensive survey of LLM-based multi-agent systems. Specifically, Section 2 introduces the core concepts, including agent environments and profiles, inter-agent interaction mechanisms, and agent-human interactions. Section 3 provides a structured review of existing studies: multi-vehicle interaction, vehicle-infrastructure interaction, and vehicle-assistant interaction. As agent capabilities continue to grow, human-vehicle co-driving is emerging as the dominant autonomous driving paradigm, with human playing an increasingly vital role. Humans collaborate with agents by providing guidance or supervising their behavior. Therefore, we consider humans as special virtual agents and examine human-agent interactions in Section 4. Section 5 explores various applications, while Section 6 compiles a comprehensive collection of public datasets and open-source resources. Section 7 discusses existing challenges and future research directions. Finally, Section 8 concludes the study.
@@ 2.2.1 Agent Environment and Profile
Table 1: Comparison of Agent Profiling Methods.
[…]
Physical environment represents the real-world setting where driver agents gather information using various sensors, such as cameras and LiDAR, and interact with other traffic participants. However, due to the high cost of vehicles and strict regulations on public roads, collecting large amounts of data in real world is impractical. As a viable alternative, the Simulation environment provides a simulated setting constructed by humans. It can accurately model specific conditions without incurring the high costs and complexities associated with real-world data collection, allowing agents to freely test actions and strategies across a variety of scenarios Dosovitskiy et al. (2017).
@@
Table 1 summarizes the advantages and limitations of different agent profiling methods in ADSs. Specifically, within Pre-defined methods, system designers explicitly define agent profiles based on prior knowledge and the analysis of complex scenarios Chen et al. (2024a). Each agent has unique attributes and behavior patterns that can be adjusted based on the scenario. In driving environments, the objectives of ADS require the collaboration of vehicle agents, infrastructure agents, and drivers. In particular, ❶ Vehicle agents denote various types of autonomous vehicles, traveling according to preset routes and traffic rules, while communicating and collaborating with other vehicles and driver agents. ❷ Infrastructure agents, e.g., traffic lights, road condition monitors, and parking facilities, provide real-time traffic information and instructions, influencing the behavior of driver and vehicle agents. However, manually crafting such roles is labor-intensive and often brittle when scenarios shift, which has stimulated interest in automatic profile construction, either generated by LLMs or extracted from large-scale datasets. Model-generated methods create agent profiles using advanced LLMs based on the interaction context and the goals that need to be accomplished Zhou et al. (2024c) and Data-derived Profile design agent profiles based on pre-existing datasets Guo et al. (2024).
@@ 2.2.2 LLM-based Multi-Agent Interaction
In LLM-based multi-agent ADS, effective communication and coordination among agents are crucial to improve collective intelligence and solve complex traffic scenarios. Agent interactions depend on both the interaction mode and the underlying interaction structure, as summarized in Table 3.
The interaction mode can be classified as: cooperative, competitive, and debate mode. ❶ In cooperative mode, agents work together to achieve shared objectives by exchanging information Chen et al. (2024d); Jin et al. (2024). ❷ In competitive mode, agents strive to accomplish their individual goals and compete with others Yao et al. (2024). ❸ The Debate mode enables agents to debate with each other, propose their own solutions, criticize the solutions of other agents, and collaboratively identify optimal strategies Liang et al. (2024).
@@
Figure 3: Different interaction modes and structures.
Specifically, ❶ the centralized interaction structures defines a central agent or a group of central agents to manage interactions among all agents Zhou et al. (2024c). ❷ The decentralized interaction structure allows for direct communication between agents, with all agents being equal to each other Hu et al. (2024b). ❸ Hierarchical structures focus on interactions within a layer or with adjacent layers Ohmer et al. (2022). ❹ The shared memory interaction structure maintains a shared message pool, allowing agents to send and extract the necessary information Jiang et al. (2024a). We provide a more detailed introduction to LLM-based multi-agent ADSs based on their interaction structures and modes in Section 3.
@@ 2.2.3 LLM-based Agent-Human Interaction
Recent studies show that human-machine co-driving systems use LLMs to improve agent-human interactions, enabling vehicles to communicate and collaborate seamlessly with human drivers through natural language Feng et al. (2024); Zou et al. (2025a, b). This allows vehicles to better understand and respond to human intent, provide context-aware responses, enhance driving safety and comfort, and offer personalized recommendations based on driver preferences. Humans also play a crucial role in guiding and supervising agent behavior, enhancing the agents’ capabilities while ensuring safety. We examine the role of humans as special virtual agents and explore agent-human interaction dynamics in Section 4.
@@ 3 LLM-based Multi-Agent Interaction
LLM-based Multi-Agent ADS Applications Collaborative Assistance-Tools ChatSim Wei et al. (2024), HumanSim Zhou et al. (2024a), EC-Drive Chen et al. (2024a), CAV-LLM-Driving-Assistant Tang et al. (2024), LLM-Edge-Device-Driving Huang et al. (2024b), ALGPT Zhou et al. (2024c) Collaborative Decision-Making LanguageMPC Sha et al. (2023), CoMAL Yao et al. (2024), AgentsCoDriver Hu et al. (2024a), AgentsCoMerge Hu et al. (2024b), CoDrivingLLM Fang et al. (2024), KoMA Jiang et al. (2024a),SurrealDriver Jin et al. (2024) Collaborative Perception Complement-Vehicle’s-FOV Dona et al. (2024), V-HOI MLCR Zhang et al. (2024a) Human-Agent Interaction Partnership Paradigm Drive-as-You-Speak Cui et al. (2024a), Reason-and-React Cui et al. (2024b), DriVLMe Huang et al. (2024a), AccidentGPT Wang et al. (2024b), ConnectGPT Tong and Solmaz (2024) Instructor Paradigm Co-Pilot Wang et al. (2023), PPE Ma et al. (2024) Multi-Agent Interaction Vehicle-Assistant Interaction ChatSim Wei et al. (2024), ALGPT Zhou et al. (2024c), AD-H Zhang et al. (2024c), SurrealDriver Jin et al. (2024), LDPD Liu et al. (2024a), V-HOI MLCR Zhang et al. (2024a) Vehicle-Infrastructure Interaction CAV-LLM-Driving-Assistant Tang et al. (2024), EC-Drive Chen et al. (2024a) Multi-Vehicle Interaction LanguageMPC Sha et al. (2023), AgentsCoDriver Hu et al. (2024a), KoMA Jiang et al. (2024a), AgentsCoMerge Hu et al. (2024b), CoDrivingLLM Fang et al. (2024),CoMAL Yao et al. (2024), Complement-Vehicle’s-FOV Dona et al. (2024)
@@ 3.1 Multi-Vehicle Interaction
Multi-vehicle interactions involve multiple autonomous vehicles powered by LLMs exchanging real-time information, such as locations, speeds, sensor data, and intended trajectories. By sharing partial observations of the environment or negotiating maneuvers, multiple vehicles overcome the inherent limitations of single-agent ADS, such as restricted perception and lack of collaboration.
Typically, these interactions operate in a cooperative mode with varying architectures. LanguageMPC Sha et al. (2023) employs a centralized structure, where a central agent acts as the fleet’s "brain," providing optimized coordination and control commands to each vehicle agent. In contrast, other decentralized methods Fang et al. (2024); Dona et al. (2024) treat all agents as peers, allowing direct vehicle-to-vehicle communication without central bottlenecks. For instance, AgentsCoDriver Hu et al. (2024a) designs an adaptive communication module that generates context-aware messages for inter-agent communication when the agent deems it necessary. AgentsCoMerge Hu et al. (2024b) and CoDrivingLLM Fang et al. (2024) incorporate agent communication directly into the reasoning process, facilitating real-time intention sharing and proactive negotiation before decision-making. Additionally, KoMA Jiang et al. (2024a) and CoMAL Yao et al. (2024) build a distributed shared memory pool, allowing agents to send and retrieve the necessary information to facilitate scalable interaction between agents.